OpenStudy (anonymous):

Solve the equation: −4(x + 10) − 6 = −3(x − 2)

OpenStudy (mimi_x3):

-2x - 40 - 6 = -3x + 6 -2x -46 = -3x + 6 x = 52

OpenStudy (anonymous):

When ever you see ( ) it means to distribute.. that means you multiply -4 by what ever is inside of ( ). so you'll get -4x - 40-6= -3x+6 then you'll have to combine ''like terms'' such as for example: 4x and -3x and whole numbers.. And you also have to try to get (x) on one side of the equal sign so you'll then have to add 3x to -4x -4x - 40-6= -3x+6 And the reason you add 3x to -4x is to cancel out the 3x +3x +3x Then you'll get -1x-40-6= 6 Then you combine -40 and -6 and get -46 so you get -1x-46=6 So then you repeat the same step: you add 46 to 6 Then you'll have -1x=52 x=-52 -Hope this helped its better to have an explanation than just the answer.